OpenStudy (curry):

ABC is a triangle with sides AB = 6m, BC =8m, and AC = 10m. A line k in the plane of the triangle ABC moves along the segment AC at the rate of 1cm per sec. The line starts at A and ends at C and is always perpendicular to AC. 1. How long does it take the line to reach the point B? 2. How long does it take the line to bisect the area of the triangle? 3. What is the area of the region that the line sweeps in its movement after 6min 45 sec? 4.What is the area of the region that the line sweeps at any give moment from the start of its movement?

OpenStudy (curry):

i solved the first one and i got 6 minutes

OpenStudy (anonymous):

ya that one i got also 6

OpenStudy (curry):

number 2 is really killing me

OpenStudy (anonymous):

sorry i suck at geometry. I dont even remeber learning this stuff in hs

OpenStudy (curry):

this is part of precalc trust me it looks simple but it is hard

OpenStudy (anonymous):

lol but it really falls under geometry i think

OpenStudy (curry):

lol where are all the genuises

OpenStudy (anonymous):

Hey sal

OpenStudy (anonymous):

I dont fall under the category of genius :D

OpenStudy (anonymous):

hey rld, haha me neither... and what a weird question. what course is this for?

OpenStudy (anonymous):

pre calc

OpenStudy (curry):

precalc

OpenStudy (anonymous):

ok so first thing is, did you do any work?

OpenStudy (curry):

its to make u use skills that u have already acquired through critical thinking

OpenStudy (anonymous):

cool never took such a course

OpenStudy (curry):

wat do u mean work i solved the first part but no matterf what i do i cant solve 2

OpenStudy (anonymous):

ok how did you get 6 minutes?

OpenStudy (anonymous):

lol ya cuz what is the height of the triangle

OpenStudy (anonymous):

ok so it is a right angle triangle, you can tell from the dimension. 6, 8, 10

OpenStudy (anonymous):

line K, is always perpendicular to AC which is the HYP. of the right angle triangle and is 10 meters.

OpenStudy (anonymous):

sal to the rescue

OpenStudy (curry):

i need 2 i found 3

OpenStudy (anonymous):

ok, you sure part 1 and 3 are right?

OpenStudy (curry):

well check my work for 3 i got 41/84 as the area

OpenStudy (curry):

41.84 as the area

OpenStudy (curry):

not 41/84

OpenStudy (curry):

not 41/84

OpenStudy (anonymous):

yeh 1 is correct.. it is 6 minutes

OpenStudy (curry):

3?

OpenStudy (curry):

wait nvm

OpenStudy (curry):

its 10.7241 as the area check that answer

OpenStudy (anonymous):

ok so for 3, you have 6*60+45=405seconds, and times cm per second you get 405 cm convert to meters. 4.05 meters, (6^2-4.05^2)^1/2= you get the third side= 4.427 area of triangle= 0.5*4.426*4.05=8.96

OpenStudy (curry):

thats wrong cause its a trap after 3.96 so if its 4.05 its a trapu cant use traingle formua

OpenStudy (curry):

trapezoiod

OpenStudy (anonymous):

ohhh haha, really is it like that?

OpenStudy (anonymous):

totally missed that

OpenStudy (curry):

ye

OpenStudy (anonymous):

then for 2, i think you have to relate the triangle at the top, with the one at the bottom interms of areas, and in terms of lengths and solve the 2 equations with 2 unknowns

OpenStudy (anonymous):

iight, i g2g hopefully you can figure it out on your own, if not maybe other ppl will be able to help but i gotta go. sorry